### v1

@@ 3.3 A bound for $J(P)$ using the graph structure of network
$\frac{1}{4(\delta_{\mathrm{in}}-1)}\mathcal{R}_{uv}\left(\mathcal{G}(P)\right)\leq\mathcal{R}_{uv}\left(\mathcal{G}(P^{*}P)\right),$
where $\delta_{\mathrm{in}}$ is the maximum indegree of the graph nodes of $\mathcal{G}_{\mathrm{dir}}(P)$.
Proof 3.18.
Let $\mathcal{G}(P)=(V,\mathcal{E})$ and $\mathcal{G}(P^{*}P)=(V,\mathcal{E}^{*})$. Observe that all the edges in $\mathcal{E}$ are also present in $\mathcal{E}^{*}$. We refer to the edges that are in $\mathcal{E}^{*}$ but not in $\mathcal{E}$ as ‘new edges’ (see Figure 2). Consider replacing each new edge in $\mathcal{G}(P^{*}P)$ with a series of two edges and an intermediate node, resulting in a modified graph $\bar{\mathcal{G}}=(\bar{V},\bar{\mathcal{E}})$. In this transformation, the resistances of the new edges are doubled. By Rayleigh’s monotonicity law, the resistance of every pair of two nodes in $V$ becomes equal to or greater than the original value and at most twice the original value, that is, $R_{uv}(\mathcal{G}(P^{*}P))\leq R_{uv}(\bar{\mathcal{G}})\leq 2R_{uv}(\mathcal{G}(P^{*}P))$ holds for every pair of nodes $u,v\in V$.
@@
Next, focus on an intermediate node $x_{uv}$ added when replacing a new edge $\{u,v\}$. For every new edge $\{u,v\}$ in $\mathcal{G}(P^{*}P)$, there exists a back-and-forth path in $\mathcal{G}(P)$. We then short-circuit the connection between $x_{uv}$ in $\bar{\mathcal{G}}$ and the pivot of the back-and-forth path. To short-circuit a connection means to reduce the resistance of a pair of nodes from infinity to zero. Applying Rayleigh’s monotonicity law again, we obtain $R_{\bar{u}\bar{v}}(\mathcal{G}^{\prime})\leq R_{\bar{u}\bar{v}}(\bar{\mathcal{G}})$ for every pair of nodes $\bar{u},\bar{v}\in\bar{V}$, where $\mathcal{G}^{\prime}$ is the resulting graph after shortening (see Fig. 3).
The graph $\mathcal{G}^{\prime}$ differs from $\mathcal{G}(P)$ only in the multiplicity of the edges, so we compare edges of both graphs. Notice that $\mathcal{G}^{\prime}$ has one self loop for each node, which is the same as $\mathcal{G}(P)$. The number of edges between $u,v\in V\ (u\neq v)$ in $\mathcal{G}^{\prime}$ is bounded by the number of back-and-forth paths through $\{u,v\}$ in $\mathcal{G}(P)$. Let $\eta$ be the upper bound of the number of such paths. Because an $n$-multiple edge is equivalent to a single edge of resistance $\frac{1}{n}$, we can use Rayleigh’s monotonicity law for $\mathcal{G}^{\prime}$ and obtain $R_{uv}(\mathcal{G}^{\prime})\geq\frac{1}{\eta}R_{uv}(\mathcal{G}(P))$. Combining the inequalities, we have
$\frac{1}{\eta}R_{uv}(\mathcal{G}(P))\leq R_{uv}(\mathcal{G}^{\prime})\leq R_{uv}(\bar{\mathcal{G}})\leq 2R_{uv}(\mathcal{G}(P^{*}P))$
for all $u,v\in\mathcal{V}$.
Figure 3: Graphs used to evaluate ¯$\bar{\mathcal{R}}(\mathcal{G}(P^{*}P))$.
@@
Finally, we evaluate $\eta$. For a back-and-forth path through $\{u,v\}$ in $\mathcal{G}(P)$, the pivot is $u$ or $v$. If $u$ is the pivot, then the possible choices of the other end of the path except $v$ are at most $\delta_{\mathrm{in}}-1$. The same is true for $v$. Therefore, we obtain $\eta\leq 2(\delta_{\mathrm{in}}-1)$ and $\frac{1}{2(\delta_{\mathrm{in}}-1)}R_{uv}(\mathcal{G}(P))\leq 2R_{uv}(\mathcal{G}(P^{*}P))$, which proves the lemma.
We are now in a position to prove Theorem 3.12.
Proof 3.19 (Proof of Theorem 3.12).
Applying Lemmas 3.15 and 3.17 to (18), we obtain
$\frac{1}{4n\pi_{\max}(\delta_{\mathrm{in}}-1)(\delta_{\mathrm{in}}+1)p_{\max}^{2}}\bar{\mathcal{R}}(\mathcal{G}(P^{*}P))\leq\bar{\mathcal{R}}(C_{P^{*}P})\leq\frac{1}{n\pi_{\min}p_{\min}^{2}}\bar{\mathcal{R}}(\mathcal{G}(P^{*}P)).$
Using Theorem 3.5 and $(\delta_{\mathrm{in}}+1)(\delta_{\mathrm{in}}-1)\leq\delta_{\mathrm{in}}^{2}$, the proof is completed.
4 Application and numerical experiment

### v2

@@ 3.3 A bound for $J(P)$ using the graph structure of network
$\frac{1}{4\delta_{\mathrm{out}}-2}\mathcal{R}_{uv}\left(\mathcal{G}(P)\right)\leq\mathcal{R}_{uv}\left(\mathcal{G}(P^{*}P)\right)\leq\mathcal{R}_{uv}\left(\mathcal{G}(P)\right),$
where $\delta_{\mathrm{out}}$ is the maximum outdegree (excluding self loops) of the graph nodes of $\mathcal{G}_{\mathrm{dir}}(P)$.
Proof 3.17.
The right-hand inequality is straightforward. For each edge $(u,v)$ in $\mathcal{G}_{\mathrm{dir}}(P)$, we can choose a back-and-forth path between $u$ and $v$, because there are edges $(u,u)$ (a self loop) and $(u,v)$ in $\mathcal{G}_{\mathrm{dir}}(P)$. Therefore, any edge appearing in $\mathcal{G}(P)$ also appears in $\mathcal{G}(P^{*}P)$. In fact,
$(P^{*}P)_{uv}=\frac{1}{\pi_{u}}\sum_{w}\pi_{w}P_{wu}P_{wv}\geq\frac{1}{\pi_{u}}\pi_{u}P_{uu}P_{uv}=P_{uu}P_{uv}>0.$
By Lemma 3.13, $\mathcal{R}_{uv}(\mathcal{G}(P^{*}P))\leq\mathcal{R}_{uv}(\mathcal{G}(P))$ for all $u,v$.
For the left-hand inequality, we adapt the standard argument for fuzz graphs [2, Lem. 5.5.1]. Write $\mathcal{G}(P)=(V,\mathcal{E})$ and $\mathcal{G}(P^{*}P)=(V,\mathcal{E}^{*})$, and call edges in $\mathcal{E}^{*}\setminus\mathcal{E}$ “new edges” (see Figure 2). Replace each new edge $\{u,v\}$ by a series of two edges with an intermediate node $x_{uv}$, preserving the total resistance (e.g., split $r$ into $r/2+r/2$). This yields a graph $\bar{\mathcal{G}}=(\bar{V},\bar{\mathcal{E}})$ that is electrically equivalent to $\mathcal{G}(P^{*}P)$, so $\mathcal{R}(\bar{\mathcal{G}})=\mathcal{R}(\mathcal{G}(P^{*}P))$ (see Figure 3).
Next, for each new edge $\{u,v\}$, choose one back-and-forth path in $\mathcal{G}(P)$ with pivot $w$ and add a zero-resistance edge between $x_{uv}$ and $w$. Identifying these endpoints yields a graph $\mathcal{G}^{\prime}$ with $\mathcal{R}(\mathcal{G}^{\prime})\leq\mathcal{R}(\bar{\mathcal{G}})$ by Rayleigh’s monotonicity (see Figure 3) [^1] Intuitively, this operation can be understood as short-circuiting.. The graph $\mathcal{G}^{\prime}$ differs from $\mathcal{G}(P)$ only in the multiplicity of edges, so we compare the two by counting parallel edges. Note that $\mathcal{G}^{\prime}$ has one self-loop at each node, as does $\mathcal{G}(P)$. For $u,v\in V$ with $u\neq v$, the number of edges between $u$ and $v$ in $\mathcal{G}^{\prime}$ is bounded by the number of back-and-forth paths through $\{u,v\}$ in $\mathcal{G}(P)$. Let $\eta$ be an upper bound on the number of such paths. Since $k$ parallel unit-resistance edges are equivalent to a single edge of resistance $1/k$, Rayleigh’s monotonicity law yields $\mathcal{R}_{uv}\left(\mathcal{G}^{\prime}\right)\geq\frac{1}{\eta}\,\mathcal{R}_{uv}\left(\mathcal{G}(P)\right)$. Combining this with the previous inequality, we obtain
$\frac{1}{\eta}\,\mathcal{R}_{uv}\left(\mathcal{G}(P)\right)\leq\mathcal{R}_{uv}\left(\mathcal{G}^{\prime}\right)\leq\mathcal{R}_{uv}\left(\bar{\mathcal{G}}\right)\leq 2\mathcal{R}_{uv}\left(\mathcal{G}(P^{*}P)\right)$
for all $u,v\in V$.
@@
Figure 3: Graphs used to evaluate $\bar{\mathcal{R}}(\mathcal{G}(P^{*}P))$.
@@
Finally, we bound $\eta$. For a back-and-forth path through $\{u,v\}$ in $\mathcal{G}(P)$, the pivot is either $u$ or $v$. If the pivot is $u$, then the other endpoint (distinct from $v$) can be chosen among at most $\delta_{\mathrm{out}}-1$ out-neighbors of $u$; the same bound holds when the pivot is $v$. Including the original edge $\{u,v\}$ of $\bar{\mathcal{G}}$, we obtain $\eta\leq 2(\delta_{\mathrm{out}}-1)+1=2\delta_{\mathrm{out}}-1$, and hence $\frac{1}{2\delta_{\mathrm{out}}-1}\,\mathcal{R}_{uv}\left(\mathcal{G}(P)\right)\ \leq\ 2\,\mathcal{R}_{uv}\left(\mathcal{G}(P^{*}P)\right)$, which proves the desired lower bound for $\bar{\mathcal{R}}\!\left(\mathcal{G}(P^{*}P)\right)$.
Applying Lemma 3.16 to $P^{*}$, we obtain the following corollary.
Corollary 3.18.
If $P^{*}P=PP^{*}$, then
$\frac{1}{4\delta_{\mathrm{in}}-2}\mathcal{R}_{uv}\left(\mathcal{G}(P)\right)\leq\mathcal{R}_{uv}\left(\mathcal{G}(P^{*}P)\right)\leq\mathcal{R}_{uv}\left(\mathcal{G}(P)\right),$
under the same assumption as Lemma 3.16.
Proof 3.19.
Substituting $P$ in Lemma 3.16 to $P^{*}$, then we obtain
$\frac{1}{4\delta_{\mathrm{in}}-2}\mathcal{R}_{uv}\left(\mathcal{G}(P^{*})\right)\leq\mathcal{R}_{uv}\left(\mathcal{G}(PP^{*})\right)\leq\mathcal{R}_{uv}\left(\mathcal{G}(P^{*})\right),$
since $(P^{*})^{*}=P$ and the outdegrees of $\mathcal{G}_{\mathrm{dir}}(P^{*})$ corresponds to the indegrees of $\mathcal{G}_{\mathrm{dir}}(P)$. By $P^{*}P=PP^{*}$ and $\mathcal{G}(P^{*})=\mathcal{G}(P)$, the proof is completed.
We are now in a position to prove Lemma 3.12.
Proof 3.20 (Proof of Lemma 3.12).
Combining Lemma 3.14 and Lemma 3.16, we obtain
$\frac{1}{n\pi_{\max}(4\delta_{\mathrm{out}}-2)(\delta_{\mathrm{in}}+1)p_{\max}^{2}}\bar{\mathcal{R}}(\mathcal{G}(P^{*}P))\leq\bar{\mathcal{R}}(C_{P^{*}P})\leq\frac{1}{n\pi_{\min}p_{\min}^{2}}\bar{\mathcal{R}}(\mathcal{G}(P^{*}P)).$
In addition, if $P^{*}P=PP^{*}$, we can change $\delta_{\mathrm{out}}$ in the leftmost term to $\delta_{\mathrm{in}}$ by Corollary 3.18. In this case, the lower bound for $\bar{\mathcal{R}}(C_{P^{*}P})$ can be obtained by $\frac{1}{n\pi_{\max}f(\delta_{\mathrm{in}})p_{\max}^{2}}\bar{\mathcal{R}}(\mathcal{G}(P^{*}P))$ where $f(\delta)=4\delta_{\mathrm{in}}^{2}+2\delta_{\mathrm{in}}-2$. Using Theorem 3.5, the proof is completed.
Theorem 3.11 follows immediately from Theorem 3.5 and Lemma 3.12.
4 Important examples and applications
In this section, we will consider important examples for applying the result in Section 3.
4.1 Necessity of the assumption $P^{*}P=PP^{*}$
We have assumed the equality $P^{*}P=PP^{*}$ for the lower bounds in Theorem 3.5 and Theorem 3.11. We will discuss the necessity of the assumption. Consider the following consensus matrix:
We enforce $\varepsilon>0$ to comply with our standing assumption that the diagonal entries are positive; otherwise some of the lemmas used in the paper do not apply.
We compute $J(P_{\varepsilon})$ and the corresponding upper and lower bounds from Theorems 3.5 and 3.11 for $\varepsilon\leq 1/2$. The results are shown in Figure 4. In Figure 4, the lower bound from Theorem 3.5 fails to hold when $\varepsilon<1/2$. This is unsurprising, since $P_{\varepsilon}^{*}P_{\varepsilon}\neq P_{\varepsilon}P_{\varepsilon}^{*}$ for $\varepsilon<1/2$, and the proof of Theorem 3.5 relies on the commutation $P^{*}P=PP^{*}$.
To understand the mechanism, note that, as $\varepsilon\to 0$, $P_{\varepsilon}^{t}-\bm{1}\bm{\pi}$ converges exponentially at a rate determined by the second largest eigenvalue of $P$. The eigenvalues of $P_{\varepsilon}$ are $1$ and $\alpha_{\pm}:=-\left(\frac{1}{4}-\varepsilon\right)\pm\frac{\mathrm{i}}{2}\sqrt{\frac{7}{4}-2\varepsilon}$. The eigenvalues of $P_{\varepsilon}^{*}$ coincide with those of $P_{\varepsilon}$ by similarity. Therefore, $(P_{\varepsilon}^{*})^{t}P_{\varepsilon}^{t}-\bm{1}\bm{\pi}$ converges exponentially to some real matrix at a rate close to $1/2$. On the other hand, $P^{*}P$ has the second eigenvalue $1-O(\varepsilon)$. This means that if $\varepsilon$ is small enough, the convergence of $(P^{*}P)^{t}$ is very slow, and its trace can be much larger than $\operatorname{tr}(P_{\varepsilon}^{*})^{t}P_{\varepsilon}^{t}$. Therefore, the right-hand side of (16) becomes much larger than the leftmost hand, and Theorem 3.5 does not hold.
However, in Figure 4, it seems that the lower bound of Theorem 3.11 holds for small $\varepsilon$, although there are no theoretical guarantees in our paper. This is likely due to the rough evaluation of Lemma 3.12. When $\varepsilon$ decreases, the average effective resistance $\bar{\mathcal{R}}(C_{P^{*}P})$ increases linearly to $1/\varepsilon$, so the evaluation of Lemma 3.12 becomes rough. Taking into account this behavior, it may be possible to find sharper or more general lower bounds for $J(P)$ in future research.
Figure 4: The values of $J(P_{\varepsilon})$ for $\varepsilon\in[0.001,0.5]$. The left panel shows the lower and upper bounds in Theorem 3.5. The right panel shows the lower and upper bounds in Theorem 3.11.
4.2 Applications to Cayley graphs
An important application where the property $P^{*}P=PP^{*}$ holds is that of Cayley consensus matrices. Given a finite abelian group $(V,+)$ and a subset $S\subseteq V$, the Cayley graph $\mathcal{G}$ is defined as a directed graph with the vertex set $V$ and the edge set $E=\{(u,v)\mid u-v\in S\}$. The associated Cayley matrix $P\in\mathbb{R}^{V\times V}$ on this graph is defined by $P_{uv}=g(u-v)$ for some function $g$, called a generator, where rows and columns are indexed by the elements of $V$. If a consensus matrix $P$ is a Cayley matrix, then $P^{*}P=PP^{*}$ holds, since Cayley matrices are normal (see [5, Section III]) and Proposition 3.3 applies.
Figure 5: Consensus algorithm on a circle.
A Cayley consensus matrix naturally arises for consensus dynamics on periodic grids (discrete tori). We begin with the one-dimensional case. As shown in Figure 5, there are $n$ agents on a circle, and each agent updates its information by the weighted sum of the left agent (with weight $p$), the right agent (with $q$) and itself (with weight 1pq1-p-q). This graph is the Cayley graph generated by the cyclic group $\mathbb{Z}_{n}:=\{0,1,\ldots,n-1\}$ with addition modulo $n$ and the generating set $S=\{-1,0,1\}$. The consensus matrix is
$P=\begin{pmatrix}1-p-q&q&0&\cdots&p\\ p&1-p-q&q&\cdots&0\\ 0&p&1-p-q&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ q&0&0&\cdots&1-p-q\\ \end{pmatrix}$.
This is a Cayley matrix obtained by
$g(x):=\begin{cases}p&\text{if }x\equiv 1\mod n\\ 1-p-q&\text{if }x\equiv 0\mod n\\ q&\text{if }x\equiv-1\mod n\\ 0&\text{otherwise}.\end{cases}$
Higher-dimensional tori are treated similarly using $\mathbb{Z}_{n}^{d}$. If we define $S$ as the neighbor set (for example, $\{-1,0,1\}^{d}$) and the value of $g(\bm{h})$ for each $\bm{h}\in S$ so that the sum of them is $1$, then we can consider a consensus algorithm on a $d$ dimensional grid where each agent communicates only with the neighbor agents. The example of a two dimensional torus is shown in Figure 6. In this case, we set $S=\{-1,0,1\}^{2}$ and $g(\pm 1,\pm 1)=0$. Grids with boundaries can be reduced to the torus case via an appropriate reflection argument [5].
Figure 6: A grid graph on two dimensional torus. The nodes in the leftmost column and the top row are adjacent to those in the rightmost column and in the bottom row, respectively.
We first recall the asymptotic behavior of $J(P)$ for Cayley consensus matrices.
Proposition 4.1 (Proposition 1 in [5], modified).
Let $\{P_{n}\}_{n\geq 1}$ be a Cayley consensus matrix family on the Cayley graph on $\mathbb{Z}_{n}^{d}$ with the neighbor set $\{-1,0,1\}^{d}$, associated with a fixed generator $g$. Then there exist $C_{d},C_{d}^{\prime}>0$ (depending only on $d$) such that:
  • if $d=1$:
    $C_{d}n^{d}\leq J(P_{n})\leq C_{d}^{\prime}n^{d}$
  • if $d=2$:
    $C_{d}\log(n^{d})\leq J(P_{n})\leq C_{d}^{\prime}\log(n^{d})$
  • if $d\geq 3$:
    $C_{d}\leq J(P_{n})\leq C_{d}^{\prime}$
This proposition says that, for the Cayley matrix $P\in\mathbb{R}^{V\times V}$ on the grid graph on the $d$ dimensional torus,
$J(P)=\begin{cases}\Theta(|V|)&\text{if }d=1\\ \Theta(\log|V|)&\text{if }d=2\\ \Theta(1)&\text{if }d\geq 3.\end{cases}$
Since $J(P)$ is already understood, we next relate $J(P)$ to our upper and lower bounds involving $\bar{\mathcal{R}}(\mathcal{G}(P))$.
Corollary 4.2.
Let $P$ be a normal consensus matrix with invariant measure $\bm{\pi}$, and let $\mathcal{G}(P)$ be the associated undirected graph with $P$. Then it holds that
$\frac{1}{p_{\max}^{2}f(\delta_{\mathrm{in}})}\bar{\mathcal{R}}(\mathcal{G}(P))\leq J(P)\leq\frac{1}{p_{\min}^{2}}\bar{\mathcal{R}}(\mathcal{G}(P)),$
where $f(\delta_{\mathrm{in}})=4\delta_{\mathrm{in}}^{2}+2\delta_{\mathrm{in}}-2$.
Proof 4.3.
Because a normal stochastic matrix is doubly stochastic, $\bm{\pi}=\frac{1}{n}\bm{1}$ and $\pi_{\min}=\pi_{\max}=\frac{1}{n}$. In addition, by the normality of $P$, $P^{*}=P$ holds and we have $P^{*}P=PP^{*}$. Applying them to Theorem 3.11, we complete the proof.
Corollary 4.2 establishes a direct relationship between the LQ cost $J(P)$ and the average effective resistance $\bar{\mathcal{R}}(\mathcal{G}(P))$ for normal consensus matrices. To further clarify the behavior of $\bar{\mathcal{R}}(\mathcal{G}(P))$ in concrete settings, we next specialize to Cayley graphs on $\mathbb{Z}_{n}^{d}$ and derive its asymptotic scaling.
Corollary 4.4.
Let $\mathcal{G}_{n}$ be the Cayley graph on $\mathbb{Z}_{n}^{d}$ with the neighbor set $\{-1,0,1\}^{d}$. Then there exist $C_{d},C_{d}^{\prime}>0$ (depending only on $d$) such that:
  • if $d=1$:
    $C_{d}n^{d}\leq\bar{\mathcal{R}}(\mathcal{G})\leq C_{d}^{\prime}n^{d}$
  • if $d=2$:
    $C_{d}\log(n^{d})\leq\bar{\mathcal{R}}(\mathcal{G})\leq C_{d}^{\prime}\log(n^{d})$
  • if $d\geq 3$:
    $C_{d}\leq\bar{\mathcal{R}}(\mathcal{G})\leq C_{d}^{\prime}$
Proof 4.5.
Let $\{P_{n}\}_{n\geq 1}$ be a Cayley consensus matrix family on $\mathcal{G}_{n}$, associated with a fixed generator $g$. Then, $p_{\min},p_{\max}$ and an upper bound of $f(\delta_{\mathrm{in}})$ is determined by $g$. By Corollary 4.2, the upper and lower bounds of $J(P_{n})$ is linear to $\bar{\mathcal{R}}(\mathcal{G}_{n})$ under fixed $g$ and $d$. Therefore, substituting $J(P_{n})$ by $\bar{\mathcal{R}}(\mathcal{G}_{n})$ in Proposition 4.1, the proof is completed.
We compare $J(P)$ and $\bar{\mathcal{R}}(\mathcal{G}(P))$ as follows:
  • We fix the dimension $d=2$ or $3$ and the size of the grid $n$. Notice that we will generate the graph with $n^{d}$ nodes.
  • For each $d$, we define the generator $f$ in the following two patterns:
    • Case 1. We set $g(\bm{h})\in(0,1]$ for each $\bm{h}\in\{-1,0,1\}^{d}$ randomly and normalize so that the sum of $g(\bm{h})$ is $1$. Then, we check whether $g(\bm{h})$ is in a certain range $p_{\min},p_{\max}]$. Here we define $p_{\min}=0.05$ and $p_{\max}=0.2$ in the case $d=2$, and $p_{\min}=0.01$ and $p_{\max}=0.1$ in the case $d=3$. In this case, the maximum indegree $\delta_{\mathrm{in}}$ of the graph is $8$ (in $d=2$) or $26$ (in $d=3$).
    • Case 2. We set $g(\bm{h})=\frac{1}{d+1}$ for each $\bm{h}\in\{\bm{0},\mathbf{e}_{1},\dots,\mathbf{e}_{d}\}$. In this case, the maximum indegree $\delta_{\mathrm{in}}$ of the graph is $2$ (in $d=2$) or $3$ (in $d=3$).
  • Generate the Cayley graph $\mathcal{G}$ and the Cayley matrix $P\in^{n^{d}\times n^{d}}$ by the generator $g$. In case 1, $P$ is randomly determined, so we generate $20$ instances of $g$ for each $d,n$.
  • Calculate $\bar{\mathcal{R}}(\mathcal{G})$ and $J(P)$. Compare the lower and upper bounds of Corollary 4.2 and $J(P)$.
Figure 7: The values of $J(P)$ in $d=2$, where $g$ is generated by the case 1. The left and right panels show the theoretical lower bound in Corollary 4.2. The right panel also shows the upper bound, which is too large to draw in the left.
Figure 8: The values of $J(P)$ in $d=3$, where $g$ is generated by the case 1. The left and right panels show the theoretical lower bound in Corollary 4.2. The right panel also shows the upper bound.
Figure 9: The values of $J(P)$ in $d=2,3$, where $g$ is generated by the case 2. The left and right panels show the theoretical lower and upper bounds in Corollary 4.2.
The result is shown in Figure 7, 8, and 9. In these figures, $J(P)$ appears to increase logarithmically in $d=2$, and does not change much in $d=3$. In addition, the change of $\bar{\mathcal{R}}(\mathcal{G})$ seems to be consistent with Corollary 4.4. These results also imply that the bound we show in the paper is not sharp in some cases. In fact, the ratio of the lower and upper bounds in the case 1 of $d=3$ is $f(26)\times 0.1^{2}/0.01^{2}=275400$. However, in the case 2 of $d=2$, all elements are the same value, so the ratio is $f(2)=18$. Thus, our bounds are most informative when the in-degree is small and the generator weights are nearly uniform.
4.3 Application to geometric graphs